OpenStudy (anonymous):

Which of the following systems of equations has infinitely many solutions? a. 4x − 7y = 1; 14y = 8x − 3 b. 8x − 3 = 6y; 12y − 6 = 2x c. 4x − 5y = 13; 10y = 8x − 26 d. 5y = -x − 5; -2x + 15y = 10

OpenStudy (anonymous):

it is the one where they both describe the same line... to find out write out each line in one form, for example slope-intercept form... the one that are the same is the choice you want.

OpenStudy (anonymous):

a) No Solution b) x=6/7 y=9/14 c) \[x = (5y +13) / 4\] and \[y \in \mathbb{R}\] d) x = -5 and y = 0

OpenStudy (anonymous):

c) And if you put x in the other equation 10y = 8(5y+13/4)-26